### v2

# Enhancement of Weak Interactions in Phase Transitions

Abstract
Weak interactions cause small parity-violating energy differences between left- and right-handed chiral systems. Although normally tiny, these effects may be significantly enhanced during collective phenomena such as phase transitions. We propose a theoretical model describing the enhancement of weak interactions in phase transitions. The enhancement factor is proportional to the critical number of atoms, $N_{c}$, in the nucleus of the new phase. After the nucleus reaches its critical size, it grows until it fills the entire system. Measurement of the ratio of produced left and right chiral structures may provide a way to measure this critical number $N_{c}$. Experiments where definite spin-chiral structures are formed during a phase transition in crossed electric and magnetic fields, indicate $N_{c}\sim 10^{9}-10^{10}$. An open question is whether a similar enhancement could operate during cosmological phase transitions - thereby boosting CP-violating effects sufficiently to contribute to the observed baryon-to-photon ratio.
I Introduction
As known, the Standard Model gives a baryon-to-photon ratio several orders of magnitude smaller than observations. In this situation it is important to check that we have not overlooked any enhancement mechanism which can play a role in electroweak baryogenesis. Indeed, systems are very sensitive to weak fields during phase transitions. A very interesting experiment was performed in Ref. [1]. One may interpret its result as a nine-orders-of-magnitude enhancement of the effects of small perturbations in phase transitions in condensed matter (see below). If a similar enhancement of CP violation existed in the electroweak phase transition (which presumably created the baryon asymmetry of the Universe), CP violation in the Standard Model might produce a baryon-to-photon ratio compatible with observations. Even if the enhancement is not large enough in the Standard Model, it may be important for other models of baryogenesis involving a phase transition.
One may think about using enhancement of weak interaction in phase transition to search for time reversal (T) and parity (P) violating interactions which were searched for in the electric dipole moment experiments with neutrons, nuclei, atoms and molecules (see e.g. [2, 3, 4, 5]).
In nature many organic molecules prefer one chiral form over another. A striking example is that almost all DNA is right-handed, while nearly all amino acids in living organisms are left-handed (see e.g.[6]). One possible explanation of this phenomenon is the small energy difference between left- and right-handed molecules (see e.g [7, 8]). This difference originates from the parity-violating component of the weak interaction between electrons and nuclei. The weak interaction mixes $s_{1/2}$ and $p_{1/2}$ states, creating a spin-helical structure within atoms [9, 2]). The spin-orbit interaction then distinguishes between coordinate and spin structures aligned in opposite directions.
It is generally believed that the parity-violating energy difference, $\Delta E_{PV}$, is too small to affect the homochirality of life. Estimate gives [10]
$\Delta E_{PV}\sim 10^{-20}Z^{5}\eta\quad\mathrm{a.u.},$ (1)
where $Z$ is the nuclear charge of the heaviest atom in the molecule and $\eta$ is a molecular asymmetry factor. The steep $Z^{5}$ dependence results from weak ($\propto Z^{3}$) and spin-orbit ($\propto Z^{2}$) scaling, but $\eta$ is typically small. In molecules with two heavy atoms, $\eta$ can be many orders of magnitude larger than in molecules with one heavy atom [10]. In heavy molecules $\Delta E_{PV}$ may exceed $10^{-11}$ eV, see e.g. [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
Energy difference between molecules of different chirality is studied using spectroscopy methods. In principle, there are other possibilities. One may study resonance chemical reactions and enhancement of weak interactions in collisions of cold molecules, where chiral molecules are formed in collision of other molecules. Weak interaction produces energy difference $\delta E_{PV}$ in the cross section resonance position in righ-hand and left-hand molecules, which results in difference in their numbers formed [24].
In a more general case, formation of chiral molecules from non-chiral components in chemical reactions create non-equal number of right-hand and left-hand molecules with the relative difference which may exceed equilibrium value $\sim\exp{(-\delta E_{PV}/T)}$. This leads to the optical activity of molecular gas or solution.
A natural place to search for enhancement of such small effects is in collective phenomena such as phase transitions - see e.g. [1]. Systems are highly sensitive to weak perturbations near criticality. In the following we develop this idea within the Zeldovich nucleation model of first-order phase transitions presented e.g. in the book [25].
[…]
II Nucleation Model and Enhancement Factor
In the nucleation picture, fluctuations create small droplets of a new phase. Droplets smaller than a critical size $r_{cr}$ quickly disappear, while larger ones grow to fill the system. The probability of forming a nucleus of radius $r$ is [25]
$S\sim e^{-W_{min}/kT},$ (2)
where $W_{min}$ is the minimal work required to form a nucleus, which may be presented as a sum of the negative volume term, reflecting advantage of the new phase, and positive surface term: $W_{min}(r)=-B\,4\pi r^{3}/3+\alpha_{t}\,4\pi r^{2}$. This function has maximum at the critical size $r_{c}=2\alpha_{t}/B$. This looks like $W_{min}(r)$ produces a potential barrier for formation of the nucleus bigger than the critical size, with $r>r_{c}$.
Using formulas presented in the book [25], it is easy to express $W_{min}(r_{c})$ in terms of the chemical potentials per particle of the old and new phases, $\mu_{1}$ and $\mu_{2}$:
$W_{min}=\frac{1}{2}N_{c}(\mu_{1}-\mu_{2}),$ (3)
where $N_{c}$ is the number of particles in the critical nucleus.
If the new phase may exist in two nearly degenerate forms (e.g. left- and right-handed chirality), then weak interactions induce a small energy difference $\Delta E$ per particle. This results in slightly different $W_{min}^{\pm}$ for the two structures. The asymmetry in their nucleation rates can be expressed as
$P\equiv\frac{S_{+}-S_{-}}{S_{+}+S_{-}}=\frac{e^{-W_{min}^{+}/kT}-e^{-W_{min}^{-}/kT}}{e^{-W_{min}^{+}/kT}+e^{-W_{min}^{-}/kT}}$ (4)
which in the linear regime ($|W_{min}^{+}-W_{min}^{-}|\ll kT$) reduces to
$P\approx-\frac{W_{min}^{+}-W_{min}^{-}}{2kT}.$ (5)
When the phase transition proceeds from a non-chiral to a chiral phase,
$P\sim-\frac{N_{c}\Delta E}{2kT}.$ (6)
Thus, the small microscopic energy difference $\Delta E$ is enhanced by the collective factor $N_{c}$, the number of particles in the critical size nucleus, which may be very large.
Here we see the potential for large enhancement: normally in statistical physics energies per degree of freedom are compared to $kT$, but a critical nucleus may contain billions of degrees of freedom, and the total difference $W_{\min}^{+}-W_{\min}^{-}$ is compared with $kT$.
Note that $N_{c}\propto(\mu_{1}-\mu_{2})^{-3}$ [25]. The difference in $\mu_{1}-\mu_{2}$ appears in metastable state of overcooled system (we assume that ordered phase appears for $T<T_{c}$). Therefore, $N_{C}$ diverges as $\mu_{1}\to\mu_{2}$ for $T\to T_{C}$. Purely theoretically, if the temperature is reduced very slowly, giving sufficient time for very unprobable, exponentially suppressed formation of very large nuclei of ordered phase, the enhancement may be nearly infinite, the system always comes to the lower energy chiral state. However, this statement is challenged by the fluctuations and effects of defects.
Another way to formulate this enhancement: the transition temperature is different for transitions to states of different chirality. So, in an ideal situation, with a very slow cooling process we may achieve transition specifically to the lower energy chiral state which has higher $T_{C}$.
In higher-order phase transitions there are no metastable phases for $T<T_{C}$, and transition always happens at $\mu_{1}=\mu_{2}$. In this sense, second-order transitions can in principle provide extreme enhancement (theoretically, $N_{c}$ may be approaching the number of particles in the system or in the correlation volume).
III Magnetic Crystals with Spin-Screw Structures
In magnetic crystals the spin-screw structures play the role of molecular chirality. The energy difference between left- and right-handed spin-screw states arises from spin-orbit coupling and external fields, see, e. g., book [2].
For the spinel ZnCr2Se4 the phase transition is classified as a weak first order phase transition (i.e. close to the second order phase transition). Near its Néel temperature $T_{N}$=21 K, experiment in crossed electric field $E=2.5$ kV/cm and magnetic field $H=12$ kOe measured an asymmetry $P\simeq-0.9$ [1]. The effect is produced by interactions with magnetic field $\hat{H}_{H}$, electric field $\hat{H}_{E}$, and spin-orbit interaction $\hat{H}_{\mathrm{SO}}$ [1, 2]. A rough estimate yields
$\Delta E\sim\frac{\langle n|\hat{H}_{\mathrm{SO}}|m\rangle\langle m|\hat{H}_{E}|k\rangle\langle k|\hat{H}_{H}|n\rangle}{(E_{n}^{(0)}-E_{m}^{(0)})(E_{n}^{(0)}-E_{k}^{(0)})}$ (7)
$\sim\frac{\,Z^{2}\alpha^{2}(e^{2}/a_{B})\cdot ea_{B}E\cdot\mu_{B}H}{(e^{2}/a_{B})^{2}}\sim 10^{-12}\mathrm{eV}\approx 10^{-8}\mathrm{K}.$ (8)
Geometric suppression factor may reduce this estimate. Using Eq. (6) we obtain
$N_{c}\gtrsim 10^{9}.$ (9)
This illustrates the potential scale of enhancement.
IV Possible implications for baryogenesis
Several studies have concluded that the Standard Model predicts a baryon-to-photon ratio eight to ten orders of magnitude smaller than observed. It is therefore crucial to investigate whether enhancement mechanisms during phase transitions could amplify CP-violating effects to observable level.
Experiments in condensed matter [1] demonstrate that tiny perturbations can strongly bias phase transitions. In ZnCr2Se4, a minuscule third-order perturbation energy shift of order $\Delta E\sim 10^{-12}$ eV $\simeq 10^{-8}$ K was sufficient to determine the chirality of spin helices in about $95\%$ of cases, at a transition temperature of $T_{N}\approx 21$ K. This corresponds to an enhancement factor of about $10^{9}-10^{10}$.
If similar enhancement of CP-violating interactions occurred during the electroweak phase transition after the Big Bang, then even the small CP violation of the Standard Model could, in principle, generate the observed baryon asymmetry. Even if insufficient, such enhancement could play an important role in extensions of the Standard Model baryogenesis scenarios.
In a static case, the energy difference between matter and antimatter requires CPT violation. However, CP violation may be sufficient in the non-stationary background of evolving Universe.
For example, in electroweak baryogenesis, bubble walls separate the high-temperature symmetric phase from the low-temperature broken-symmetry phase. CP-violating interactions modify the transmission coefficients $T_{\pm}$ for particles and antiparticles crossing these walls (see, e.g., Refs. [28, 29, 30, 31]). According to Ref. [29], the relative difference in the transmission coefficients for quarks and antiquarks is $\Delta\sim 10^{-4}$ in the Standard Model, while the results of Refs. [30, 31] are significantly smaller. This effect leads to different numbers of particles and antiparticles inside the bubble; in the symmetric phase outside the bubble, rapid equilibration occurs due to sphaleron processes.
The resulting differences in reflection and transmission affect both the internal pressures and the critical bubble sizes, thereby producing an asymmetry in bubble nucleation rates. Indeed, the pressure is generated by reflection, therefore it is reduced by transmission: $P_{\pm}=(1-T_{\pm})P_{0}$, where $P_{0}$ is the pressure for $T_{\pm}=0$. The pressure difference implies a difference in the minimal work,
$d(W_{\min}^{+}-W_{\min}^{-})\sim(P_{+}-P_{-})\,dV.$ (10)
Another effect is the difference in critical sizes $r_{\pm}$ of bubbles containing particles and antiparticles. The mechanical equilibrium condition depends on the pressure inside and outside the bubble and on the radius $r$:
$P_{\mathrm{in}}+B=P_{\mathrm{out}}+\frac{2\alpha_{t}}{r},$ (11)
where $B$ is the (volume) bag constant (from the Higgs-field energy) and $\alpha_{t}$ is the surface tension. Therefore a pressure difference between particle and antiparticle cases produces a difference in the critical radius, giving an additional contribution to $W_{\min}^{+}-W_{\min}^{-}$. These two effects may lead to a “natural selection”: critical bubbles containing matter appear more frequently than those containing antimatter.
Although we can not provide here any quantitative estimates, the condensed-matter analogy suggests that enhancements by factors of order $10^{9}-10^{10}$ are plausible. Such amplification could make CP-violating effects during the electroweak phase transition a viable contributor to baryogenesis.
An open question is whether the nucleation framework and its $N_{c}$-based enhancement can be extended to a crossover rather than a sharp first-order transition. Even in a crossover, droplet-like regions (”bubbles”) of the emergent phase appear within the symmetric phase; see, e.g., [32] and references therein.
V Conclusion
We have analyzed a simple mechanism by which energy differences are amplified during phase transitions. In the Zeldovich nucleation framework the macroscopic bias
$P\sim-\frac{N_{c}\,\Delta E}{2kT}$
links a microscopic splitting $\Delta E$ to a collective factor $N_{c}$ - the number of particles in the critical nucleus of ordered phase. Enhancement factor $N_{c}$ grows rapidly as the temperature of overcooled material approaches critical transition temperature, $T\to T_{c}$, and can become extremely large in weakly first-order or effectively second-order cases.
The ZnCr2Se4 data [1] suggest an effective enhancement consistent with $N_{c}\sim 10^{9}-10^{10}$, illustrating the scale attainable in real materials.
Beyond condensed matter, the same logic may appliy to cosmological phase transitions: if analogous enhancement acts on CP-violating terms during the electroweak transition, the effective amplification may help bridge the gap between Standard-Model CP violation and the observed baryon asymmetry, or significantly impact scenarios beyond the Standard Model.
Acknowledgements.— I am grateful to Elizabeth Angstmann and Jacinda Ginges for their contributions during the early stages of this work. The work was supported by the Australian Research Council Grant No. DP230101058.